OpenStudy (anonymous):

Find an equation of the given line containing the given pair of points (-2,-5) and (-8,-4)

OpenStudy (anonymous):

slope = 1/6 ===> y=(1/6)x+b ===> b= -14/3 y = (1/6)x-14/3

OpenStudy (anonymous):

m=-1/6 then -5=-1/6(-2)+c c=-16/3 y=-1/6(x)-16/3